OpenStudy (anonymous):

The cost of a graphing calculator today is three-fourths of the cost of the calculator 5 years ago. The cost of the graphing calculator today is $63. Find the cost of the calculator 5 years ago.

OpenStudy (anonymous):

\[P(\frac{3}{4})=63\] \[P=63\times \frac{4}{3}\]

OpenStudy (lilg132):

47.25

OpenStudy (lilg132):

63 * 0.75

OpenStudy (anonymous):

i tried it

OpenStudy (anonymous):

any more

OpenStudy (akshay_budhkar):

P*3/4=63 therefore P=63*4/3 therefore P=84

OpenStudy (akshay_budhkar):

um do u get it?

OpenStudy (anonymous):

i love you lol